OpenStudy (anonymous):

Which mixed number is greater than 4.63? A. 4 2/3 B. 4 3/5 C. 4 5/9 D. 4 5/8

OpenStudy (anonymous):

@xNightwolfx

OpenStudy (anonymous):

A for sure

OpenStudy (anonymous):

again...2/3 is .6666 lol

OpenStudy (anonymous):

well what is 4.63 as a fraction?

OpenStudy (anonymous):

Well it is asking which is greater right? If you don't know how to get a decimal out of a fraction then go down that list and just do 3/5 in a caculator which is .6 which is less than .63

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so your sure?

OpenStudy (anonymous):

5/9 is .555

OpenStudy (anonymous):

Positive

OpenStudy (anonymous):

ok... if i got my last question wrong i will blame u XD

OpenStudy (anonymous):

ok phew 100%

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

:) Your welcome

OpenStudy (anonymous):

@xNightwolfx can you help with a few others maybe? i had 2 assessments

OpenStudy (anonymous):

Depends on what they are :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is the least common denominator for the fractions 9/14 and 2/21

OpenStudy (anonymous):

A. 7 B. 42 C. 84 D. 294

OpenStudy (anonymous):

LCD (Least common denominator) for the first one is 42...90% sure on that one.

OpenStudy (anonymous):

Oh yeah definitely 42

OpenStudy (anonymous):

Whats the other question?

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

what fraction is equal to 5 3/4

OpenStudy (anonymous):

A. 19/4 B. 23/4 C. 8/4 D. 15/4

OpenStudy (anonymous):

i have 3 others after this btw

OpenStudy (anonymous):

B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Btw this is bringing back to previous years of learning so I may not be 100% all the time

OpenStudy (anonymous):

what mixed number is equal to 904/100 A. 9 2/5 B. 9 4/10 C. 9 4/25 D. 9 1/25

OpenStudy (anonymous):

D

OpenStudy (anonymous):

i think its b maybe

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

Definitely D

OpenStudy (anonymous):

this one has stuff i cant like type so one sec

OpenStudy (anonymous):

number lines

OpenStudy (anonymous):

Can you screenshot it or draw it?

OpenStudy (anonymous):

Harold made the conjecture that the sum of any two prime numbers is a composite number. Which equation is a counterexample to his conjecture? A. 4 + 1 = 5 B. 3 + 10 = 13 C. 17 + 2 = 19 D. 11 + 5 = 16 this is the next one

OpenStudy (anonymous):

C

OpenStudy (anonymous):

lets see....

OpenStudy (anonymous):

80 % i passed the one i got wrong was the one with the number line

OpenStudy (anonymous):

I would've gotten you 100% if I knew what that line was ;)

OpenStudy (anonymous):

Nice job btw!

OpenStudy (anonymous):

lol ok thx i would give you another medal but i cant

OpenStudy (anonymous):

can't

OpenStudy (anonymous):

That's alright, like I said I'm here to help :)

OpenStudy (anonymous):

what gender are you btw?

OpenStudy (anonymous):

I'm a dude

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Anyways, good luck to you in your future studies! :D

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

ok thanks bye